# Expert-Token Resonance: Redefining MoE Routing through Affinity-Driven Active Selection

### v2

@@ Conclusion
In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs. Building upon the feature extraction component and locality loss of LocMoE, our approach introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates. By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized. Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC. Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, our approach outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively. Both theoretical proofs and experimental results demonstrate the benefits of the proposed routing strategy for LLM training efficiency. Future work may explore methods to compress communication data to further reduce communication overhead.
Appendix A Appendix
Appendix B Missing Proof
@@
$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$
@@
$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$
@@
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$ (12)
@@
$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
@@
$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
@@
Appendix C Token Feature Distribution
@@ Appendix D Experimental Setup
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval (Huang et al. 2024) and TeleQnA (Maatouk et al. 2023) to evaluate the model’s Chinese language capability.

### v3

@@ Conclusion
In this paper, we propose a novel expert routing framework that enhances MoE efficiency through three key innovations: an efficient routing mechanism with lightweight computation, a bidirectional expert-token resonance selection mechanism, which combined ECR and TCR, and a dynamic capacity bounds module. The framework integrates orthogonal feature extraction and optimized expert localization loss, effectively addressing expert homogeneity while improving routing performance. Our local expert strategy demonstrates advantages in both load balancing and communication efficiency. Experimental results validate the effectiveness of the proposed framework across multiple benchmarks. Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all models are evaluated with the open-source datasets GPQA and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, our approach outperforms the Baseline by 14.1%, 13.6%, and 9.7% on GDAD, GPQA, and TeleQnA, respectively. Future work may explore methods to compress communication data to further reduce communication overhead.
Appendix
Missing Proof
@@
$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$ (13)
@@
$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$ (14)
@@
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$
@@
$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$ (15)
@@
$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$ (16)
@@
Token Feature Distribution
@@
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval (Huang et al. 2024b) and TeleQnA (Maatouk et al. 2023) to evaluate the model’s Chinese language capability.